Orly's age is 4 years more than thrice Perla's age. Four years ago, Orly was four times as old as Perla was then. How old are they now?
Answer by ankor@dixie-net.com(22740)

Orly's age is 4 years more than thrice Perla's age.
o = 2p+4
:
Four years ago, Orly was four times as old as Perla was then.
o-4 = 4(p-4)
o = 4p - 16 + 4
o = 4p - 12
Replace o with 2p+4
2p + 4 = 4p - 12
4 + 12 = 4p - p
16 = 2p
p = 16/2
p = 8 yrs is Perla's age now
:
You can find Orly's age now, check solutions in the 2nd statement